# Sequence Reconstruction over the Deletion Channel

### v1

@@ 3.2. The Upper Bound
The proof follows from the proof of Theorem 8 in [3]. Let $\chi=\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap\mathcal{D}_{t}(x_{3}).$ We have
@@ 4.2. The Upper Bound
$\ell=3$ , $t\geqslant 1$, and $n\geqslant t+\ell-1$, in the previous section it has been proven.
@@
Suppose in the following vectors $x_{1}$$_{2},…,xx_{\ell}$ are such that $\ell_{1}$ of them start with $10$ and $\ell_{2}$ of them start with $01$, where $\ell_{1}\geqslant 2$ and $\ell_{2}\geqslant 2$. Then
@@
Let $\ell,\ell_{1},\ell_{2}\geqslant 2$ with $\ell_{1}+\ell_{2}=\ell$ and $\ell\geqslant 5$. We will show that
@@
The proof follows from the proof of Theorem 8 in [1]. The proof will be done by induction. The base case when $n=6$ can be verified by a computer. Suppose that the result holds for all $m<n.$ Let $\chi=\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap\cdots\cap\mathcal{D}_{t}(x_{\ell})$. We have
@@
Suppose in the following vectors $x_{1}$$_{2},…,xx_{\ell}$ are such that $\ell_{1}\geqslant 1$ of them starts with $11$ and $\ell_{3}\geqslant 1$ of them starts with $00$. Then
@@
Suppose in the following vectors $x_{1}$$_{2},…,xx_{\ell}$ are such that $\ell_{1}\geqslant 1$ of them starts with $11$, $\ell_{2}\geqslant 1$ of them starts with $01$, and $\ell_{4}$ of them starts with $10$. Then
@@
We have $D(n-3,t-2)=\sum_{i=0}^{t-2}\binom{n-t-1}{i}$.
$\chi^{1}\subset 1\circ\max_{y_{1},y_{2}\in\{0,1\}^{n-2}}\mathcal{D}_{t-1}(y_{1})\cap\mathcal{D}_{t-1}(y_{2})$.
@@
$\chi^{0}$ $\subset\max_{x_{i}\in\{0,1\}^{n-2},y\in\{0,1\}^{n-k}}\cap_{i=1}^{\ell_{4}}\mathcal{D}_{t-1}(x_{i})\cap\mathcal{D}_{t-k+1}(y)$
$\subset\max_{y\in\{0,1\}^{n-3}}\mathcal{D}_{t-2}(y).$
@@
Suppose in the following vectors $x_{1}$$_{2},…,xx_{\ell}$ are such that one of them starts with $10$ and one of them starts with $11$. For this case due to symmetry it is the same as the above lemma.

### v2

@@ 3.2. The Upper Bound
Our proof follows the proof of Theorem 8 in [3]. Let $\chi=\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap\mathcal{D}_{t}(x_{3}).$ We have
@@ 4.2. The Upper Bound
For $\ell=3$ , $t\geqslant 1$, and $n\geqslant t+\ell-1$, in the previous section it has been proven.
@@
Suppose that among the vectors $x_{1},x_{2},\dots,x_{\ell}$, $\ell_{1}$ of them begin with $10$ and $\ell_{2}$ of them begin with $01$, where $\ell_{1}\geqslant 2$ and $\ell_{2}\geqslant 2$. Then
@@
Let $\ell,\ell_{1},\ell_{2}\geqslant 2$ with $\ell_{1}+\ell_{2}=\ell$ and $\ell\geqslant 4$. We will show that
@@
Our proof follows the proof of Theorem 8 in [3]. Let $\chi=\mathcal{D}_{t}(x_{1})\cap\mathcal{D}_{t}(x_{2})\cap\cdots\cap\mathcal{D}_{t}(x_{\ell})$. We have
@@
Suppose that among the vectors $x_{1},x_{2},\dots,x_{\ell}$, $\ell_{1}\geqslant 1$ of them begin with $11$ and $\ell_{3}\geqslant 1$ of them begin with $00$. Then
@@
Suppose that among the vectors $x_{1},x_{2},\dots,x_{\ell}$, $\ell_{1}\geqslant 1$ of them begin with $11$, $\ell_{2}\geqslant 1$ of them begin with $01$, and $\ell_{4}$ of them begin with $10$. Then
@@
Similarly, $\chi^{1}\subset 1\circ\max_{y_{1},y_{2}\in\{0,1\}^{n-2}}\mathcal{D}_{t-1}(y_{1})\cap\mathcal{D}_{t-1}(y_{2})$.
@@
$\chi^{0}$ $\subset 0\circ\max_{x_{i}\in\{0,1\}^{n-2},y\in\{0,1\}^{n-k}}\cap_{i=1}^{\ell_{4}}\mathcal{D}_{t-1}(x_{i})\cap\mathcal{D}_{t-k+1}(y)$
$\subset 0\circ\max_{y\in\{0,1\}^{n-3}}\mathcal{D}_{t-2}(y).$
@@
Suppose that among the vectors $x_{1},x_{2},\dots,x_{\ell}$, one begins with $10$ and another begins with $11$. In this case, by symmetry, the situation is equivalent to that considered in the preceding lemma.